# A Search for Binary Black Hole Mergers in LIGO O1–O3 Data with Convolutional Neural Networks

### v1

@@ Abstract
Since the first detection of gravitational waves in 2015 by LIGO from the binary black hole merger GW150914, gravitational wave astronomy has developed significantly, with over 200 compact binary merger events cataloged. The use of neural networks has the potential to significantly speed up the detection, classification, and especially parameter estimation for gravitational wave events, compared to current techniques, quite important for electromagnetic follow-up of events. In this work, we present a machine learning pipeline using neural networks to detect gravitational wave events. We generate training data using real LIGO data to train and refine neural networks that can detect binary black hole (BBH) mergers, and apply these models to search through LIGO’s first three observing runs. We detect 57 out of the 75 total cataloged BBH events with two detectors of data in O1, O2, and O3, with 57 false positives that can mostly be ruled out with parameter inference and human inspection. Finally, we extensively test this pipeline on time-shifted data to characterize its False Alarm Rate (FAR). These results are an important step in developing machine learning-based GW searches, enabling low-latency detection and multi-messenger astronomy.
@@ I Introduction
Gravitational-wave (GW) astronomy has opened up a new way of observing and exploring the universe. Following the first detection of gravitational waves from a binary black hole (BBH) merger (Abbott_2016) by the two Laser Interferometer Gravitational-Wave Observatory (LIGO) detectors (Aasi_2015) in 2015, there are now over 200 compact binary coalescence (CBC) detections reported by the LIGO-Virgo-KAGRA (LVK) collaboration during the first four observing runs, which include a total of 51 months from 2015–2025. Due to the increasing sensitivity of the interferometers, the number of CBC detections has increased in each observing run, with 3 detections during O1, 8 detections during O2, 79 detections during O3, and over 200 detections during the first 23 months of the recently-completed O4 (Abbott_2023b; LIGO_2025a). While most of the detected events are BBH mergers, they also include a small number of binary neutron star (BNS) and neutron star-black hole (NSBH) mergers (Abbott_2017a; Abbott_2020; Abbott_2021b; LIGO_2025b; LIGO_2025c). With future ground- and space-based detectors, the number of detections is expected to increase dramatically, paving the way for GW astrophysics to become a primary tool in understanding compact objects and their formation pathways.
@@
The basic structure of current CBC detection algorithms is based on matched-filtering, whereby candidates are identified by comparing the data against sets of pre-calculated templates, providing a discrete sampling of the parameter space, specifically by the mass of each binary component, and their dimensionless spins (Wainstein_1962; Abbott_2020b). The matched-filter correlation provides a time-series of signal-to-noise (SNR) values, and high SNR values trigger further analysis. Events with consistent parameters across detectors are assigned a ranking statistic and are subject to further analysis, including full parameter estimation of the other parameters (orientation, sky location, etc.) (Abbott_2023b).
@@
Machine learning (ML) presents an opportunity for improving the speed, and possibly even accuracy, of GW detections and parameter estimation, enabling more rapid follow-up of any electromagnetic counterparts of future multi-messenger events (Krastev_2021; Qiu_2023). Since the first application of ML to GW detections, starting in 2018 (George_2018a), there have been several efforts to explore and enhance ML methods for the detection of CBCs. For example, George_2018a; Gabbard_2018 showed that deep learning (DL) with convolutional neural networks (CNNs) could detect simulated BBH mergers in time-series data with Gaussian noise and reach accuracies comparable to matched-filtering, as well as being more computationally efficient; this work was extended shortly thereafter using real LIGO data to detect both simulated and real BBH mergers, with similar sensitivities to matched-filtering (George_2018b). Krastev_2020 was the first to use ML to search for simulated GW signals of BNS mergers and demonstrated that a CNN could distinguish them from simulated LIGO noise and BBH mergers. In subsequent work, Krastev_2021 and Qiu_2023 extended the CNN framework to detect and estimate the parameters of GWs from BNS and NSBH mergers, respectively, using real LIGO noise.
@@
However, these early searches only evaluated ML models on time periods with known GW signals so they could not properly account for false positives. Recently, direct ML searches for BBH signals in large sections of LIGO data have been performed, allowing for proper calculation of false alarm rates. In particular, Koloniari_2025 used AresGW and Marx_2025 used AFrame to search through O3 data for BBH mergers, detecting 40 and 38 of the 65 cataloged BBH mergers in O3 with two detectors of data.
@@
In this paper we present an updated ML algorithm for detecting BBH mergers (extending our work in Krastev_2020; Krastev_2021; Qiu_2023), and carry out a full search for BBH mergers in all LIGO O1–O3 data. The paper is organized as follows. In §II we describe our methods for generating training data and training time-series and time-frequency neural network models for BBH detection. We present the results of our search on both real LIGO two-detector data, and on time-shifted data (to calculate false alarm rates) in §III. Finally, we discuss the performance of our ML pipeline, the detected and missed events, and false positives in §IV.
@@ II Methods
The overall structure of our process follows our previous work in Krastev_2020; Krastev_2021; Qiu_2023. We first generate a large template dataset of real LIGO noise with injected BBH signals. We then train a CNN on these data to classify between the signal and noise templates.
Our previous papers evaluated the neural network’s ability to correctly classify individual events using only the data centered around known detection times Krastev_2020; Krastev_2021; Qiu_2023. The events were considered “detected” if the model assigned a probability of $p>0.5$ of the correct category. However, here we perform a “blind” search of the entire LIGO data stream from O1–O3. This allows us to properly take into account false positives and to calculate false alarm rates. We consider events to be “detected” if we recover them from all of the data, with detection thresholds chosen to minimize the false alarm rate.
@@ II.1 Dataset Construction
We obtain LIGO data from the LIGO and Virgo Gravitational Wave Open Science Center (GWOSC) (Abbott_2021a; Abbott_2023a). Specifically, for the training data we use about $22$ days of continuous O2 data from the LIGO Livingston (L1) and Hanford (H1) detectors sampled at 4096 Hz, in which both detectors are operational and no known GW events or hardware injections are present. To simulate BBH signals, we use the LIGO Algorithm Library Suite (LALSuite) (lalsuite) to generate waveforms through the PyCBC Python library (Nitz_2024). In particular, we use the SEOBNRv4 time domain approximant (Bohe_2017).
@@
For all waveforms, we assume zero spin and zero eccentricity. We sample the waveforms at 4096 Hz for 4 s. After projecting the waveforms to each LIGO detector, we shift them in time to position the peak of the signal randomly in the window of 3.73.93.7-3.9 s within the 4 s template, or 3.793.813.79-3.81 s for high-resolution models (see §II.4). We whiten the noise data and the simulated signals separately with the power spectral density (PSD) computed directly from the raw GW strain data using Welch’s method (Welch_1967). We scale the amplitude of the waveforms to reach a certain sampled SNR level; sampling SNR is equivalent to sampling distance, but we find it effective to draw SNR from a uniform distribution: $\textrm{SNR}=\textrm{Unif}(7,20)$. An example of a generated BBH template is shown in Figure 2.
@@ II.3 Time-Frequency Templates
While using the 1D time series template data is useful, we find that it is beneficial to combine this with 2D time-frequency data, computed with the Q-transform available with the GWpy package (gwpy). These templates are generated in the exact way as the time-series data, but with the Q-transform computed from 30 to 330 Hz, using the last 0.5 s of the template, and generating a $300x300$ array. This can cut off the early part of some longer-duration waveforms, but we choose this size as a balance between duration, resolution, and data size. An example of this process is shown in Figure 2, which shows a 0.5 s window around the signal peak in the time-series data with its corresponding Q-transform.
@@ II.7 Neural Network Architecture and Training
For our neural network models, we use a structure similar to that of Qiu_2023 for all models. We use several different model architectures for time-series, using 1D layers, and time-frequency, using 2D layers (see Table 1). The “CNN” architecture contains 4 convolutional layers with ReLU activation, each followed by a max pooling layer, then followed by 2 dense hidden layers. The number of filters in each convolutional layer is 32, 64, 128, and 256, respectively. For the time-series models, the kernel sizes in each ResNet block are 16, 8, 8, and 8, respectively, and for the time-frequency models, the kernel sizes are all 3. The max pooling layers are size 4 for time-series and size 2 for time-frequency. The dense hidden layers have widths of 128 and 64. The final softmax output corresponds to the number of predicted classes, 2, defined as BBH and noise.
Most of our models use a ResNet structure, which involves residual connections: the output of an earlier layer in a block is added to the output of a later layer. These residual connections are designed to make deeper neural networks easier to optimize and achieve high accuracies (He_2016). In our ResNet architecture, each of the 4 convolutional layers in our CNN architecture is replaced by a ResNet blocks of 3 convolutional layers and a batch normalization layer each, as shown in the right panel of Figure 3. For one model, we add a dropout layer with a rate of 0.5 before the dense layers. As an example, the schematic diagram of the architecture of the two-detector time-series ResNet model is shown in Figure 3. It has 3,862,756 parameters in total. Finally, we point out that the exact hyperparameters we use for our models are unlikely to change the results significantly. We have chosen the specific hyperparameters to give good performance, but a thorough hyperparameter optimization would only have a very small effect on the pipeline’s results.
We build and train our neural network models using TensorFlow (Tensorflow_2015). We use the Adam optimizer (Kingma_2017) with AMSgrad (Reddi_2019). We use $\alpha=0.001$ as our learning rate, $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, and a batch size of 216. We use sparse categorical cross-entropy loss and train for 200 epochs. Our final models are taken from the epoch with the lowest validation loss. We train our models using 4 NVidia A100 GPUs with a data-parallel strategy.
@@ IV.4 Comparison to Other Searches
Now, we will perform a brief comparison of the results of our search to the two most comparable ML-based searches that have been done, with the algorithms AresGW (Koloniari_2025), and AFrame (Marx_2025). Both AresGW and AFrame only searched for BBH events in O3, while we included O1 and O2. With $p_{\textrm{astro}}>0.5$, AresGW detects 40 of the 65 total BBH events in GWTC-3 with two detectors of data and no data quality issues, while AFrame detects 38 of them. Our search detects 49 of the events, although some of those are detected with fairly high False Alarm Rates.

### v2

@@ Abstract
Since the first detection of gravitational waves in 2015 by LIGO from the binary black hole merger GW150914, gravitational-wave astronomy has developed significantly, with over 200 compact binary merger events cataloged. The use of neural networks has the potential to significantly speed up the detection, classification, and especially parameter estimation for gravitational wave events, compared to current techniques, quite important for electromagnetic follow-up of events. In this work, we present a machine learning pipeline using neural networks to detect gravitational wave events. We generate training data using real LIGO data to train and refine neural networks that can detect binary black hole (BBH) mergers, and apply these models to search through LIGO’s first three observing runs. We detect 57 out of the 75 total cataloged BBH events with two detectors of data in O1, O2, and O3, with 57 false positives that can mostly be ruled out with parameter inference and human inspection. Finally, we extensively test this pipeline on time-shifted data to characterize its False Alarm Rate (FAR). These results are an important step in developing machine learning-based GW searches, enabling low-latency detection and multi-messenger astronomy.
@@ I Introduction
Gravitational-wave (GW) astronomy has opened up a new way of observing and exploring the universe. Following the first detection of gravitational waves from a binary black hole (BBH) merger (Abbott and others, 2016) by the two Laser Interferometer Gravitational-Wave Observatory (LIGO) detectors (Aasi and others, 2015) in 2015, there are now over 200 compact binary coalescence (CBC) detections reported by the LIGO-Virgo-KAGRA (LVK) collaboration during the first four observing runs, which include a total of 51 months from 2015–2025. Due to the increasing sensitivity of the interferometers, the number of CBC detections has increased in each observing run, with 3 detections during O1, 8 detections during O2, 79 detections during O3, and over 200 detections during the first 23 months of the recently-completed O4 (Abbott and others, 2023a; Abac and others, 2025a). While most of the detected events are BBH mergers, they also include a small number of binary neutron star (BNS) and neutron star-black hole (NSBH) mergers (Abbott and others, 2017, 2020b, 2021a; Abac and others, 2025b, c). With future ground- and space-based detectors, the number of detections is expected to increase dramatically, paving the way for GW astrophysics to become a primary tool in understanding compact objects and their formation pathways.
@@
The basic structure of current CBC detection algorithms is based on matched-filtering, whereby candidates are identified by comparing the data against sets of pre-calculated templates, providing a discrete sampling of the parameter space, specifically by the mass of each binary component, and their dimensionless spins (Wainstein and Zubakov, 1962; Abbott and others, 2020a). The matched-filter correlation provides a time-series of signal-to-noise (SNR) values, and high SNR values trigger further analysis. Events with consistent parameters across detectors are assigned a ranking statistic and are subject to further analysis, including full parameter estimation of the other parameters (orientation, sky location, etc.) (Abbott and others, 2023a).
@@
Machine learning (ML) presents an opportunity for improving the speed, and possibly even accuracy, of GW detections and parameter estimation, enabling more rapid follow-up of any electromagnetic counterparts of future multi-messenger events (Krastev et al., 2021; Qiu et al., 2023). Since the first application of ML to GW detections, starting in 2018 (George and Huerta, 2018a), there have been several efforts to explore and enhance ML methods for the detection of CBCs. For example, George and Huerta (2018a); Gabbard et al. (2018) showed that deep learning (DL) with convolutional neural networks (CNNs) could detect simulated BBH mergers in time-series data with Gaussian noise and reach accuracies comparable to matched-filtering, as well as being more computationally efficient; this work was extended shortly thereafter using real LIGO data to detect both simulated and real BBH mergers, with similar sensitivities to matched-filtering (George and Huerta, 2018b). Krastev (2020) was the first to use ML to search for simulated GW signals of BNS mergers and demonstrated that a CNN could distinguish them from simulated LIGO noise and BBH mergers. In subsequent work, Krastev et al. (2021) and Qiu et al. (2023) extended the CNN framework to detect and estimate the parameters of GWs from BNS and NSBH mergers, respectively, using real LIGO noise.
@@
However, these early searches only evaluated ML models on time periods with known GW signals so they could not properly account for false positives. Recently, direct ML searches for BBH signals in large sections of LIGO data have been performed, allowing for proper calculation of false alarm rates. In particular, Koloniari et al. (2025) used AresGW and Marx et al. (2025) used AFrame to search through O3 data for BBH mergers, detecting 40 and 38 of the 65 cataloged BBH mergers in O3 with two detectors of data.
@@
In this paper we present an updated ML algorithm for detecting BBH mergers (extending our work in Krastev 2020; Krastev et al. 2021; Qiu et al. 2023), and carry out a full search for BBH mergers in all LIGO O1–O3 data. The paper is organized as follows. In §II we describe our methods for generating training data and training time-series and time-frequency neural network models for BBH detection. We present the results of our search on both real LIGO two-detector data, and on time-shifted data (to calculate false alarm rates) in §III. Finally, we discuss the performance of our ML pipeline, the detected and missed events, and false positives in §IV.
@@ II Methods
The overall structure of our process follows our previous work in Krastev (2020); Krastev et al. (2021); Qiu et al. (2023). We first generate a large template dataset of real LIGO noise with injected BBH signals. We then train a CNN on these data to classify between the signal and noise templates.
Our previous papers evaluated the neural network’s ability to correctly classify individual events using only the data centered around known detection times Krastev (2020); Krastev et al. (2021); Qiu et al. (2023). The events were considered “detected” if the model assigned a probability of $p>0.5$ of the correct category. However, here we perform a “blind” search of the entire LIGO data stream from O1–O3. This allows us to properly take into account false positives and to calculate false alarm rates. We consider events to be “detected” if we recover them from all of the data, with detection thresholds chosen to minimize the false alarm rate.
@@ II.1 Dataset Construction
We obtain LIGO data from the LIGO and Virgo Gravitational Wave Open Science Center (GWOSC) (Abbott and others, 2021b, 2023b). Specifically, for the training data we use about $22$ days of continuous O2 data from the LIGO Livingston (L1) and Hanford (H1) detectors sampled at 4096 Hz, in which both detectors are operational and no known GW events or hardware injections are present. To simulate BBH signals, we use the LIGO Algorithm Library Suite (LALSuite) (LIGO Scientific Collaboration et al., 2018) to generate waveforms through the PyCBC Python library (Nitz and others, 2024). In particular, we use the SEOBNRv4 time domain approximant (Bohé and others, 2017).
@@
For all waveforms, we assume zero spin and zero eccentricity. We sample the waveforms at 4096 Hz for 4 s. After projecting the waveforms to each LIGO detector, we shift them in time to position the peak of the signal randomly in the window of 3.73.93.7-3.9 s within the 4 s template, or 3.793.813.79-3.81 s for high-resolution models (see §II.4). We whiten the noise data and the simulated signals separately with the power spectral density (PSD) computed directly from the raw GW strain data using Welch’s method (Welch, 1967). We scale the amplitude of the waveforms to reach a certain sampled SNR level; sampling SNR is equivalent to sampling distance, but we find it effective to draw SNR from a uniform distribution: $\textrm{SNR}=\textrm{Unif}(7,20)$. An example of a generated BBH template is shown in Figure 2.
@@ II.3 Time-Frequency Templates
While using the 1D time series template data is useful, we find that it is beneficial to combine this with 2D time-frequency data, computed with the Q-transform available with the GWpy package (Macleod et al., 2021). These templates are generated in the exact way as the time-series data, but with the Q-transform computed from 30 to 330 Hz, using the last 0.5 s of the template, and generating a $300x300$ array. This can cut off the early part of some longer-duration waveforms, but we choose this size as a balance between duration, resolution, and data size. An example of this process is shown in Figure 2, which shows a 0.5 s window around the signal peak in the time-series data with its corresponding Q-transform.
@@ II.7 Neural Network Architecture and Training
For our neural network models, we use a structure similar to that of Qiu et al. (2023) for all models. We use several different model architectures for time-series, using 1D layers, and time-frequency, using 2D layers (see Table 1). The “CNN” architecture contains 4 convolutional layers with ReLU activation, each followed by a max pooling layer, then followed by 2 dense hidden layers. The number of filters in each convolutional layer is 32, 64, 128, and 256, respectively. For the time-series models, the kernel sizes in each ResNet block are 16, 8, 8, and 8, respectively, and for the time-frequency models, the kernel sizes are all 3. The max pooling layers are size 4 for time-series and size 2 for time-frequency. The dense hidden layers have widths of 128 and 64. The final softmax output corresponds to the number of predicted classes, 2, defined as BBH and noise.
Most of our models use a ResNet structure, which involves residual connections: the output of an earlier layer in a block is added to the output of a later layer. These residual connections are designed to make deeper neural networks easier to optimize and achieve high accuracies (He et al., 2016). In our ResNet architecture, each of the 4 convolutional layers in our CNN architecture is replaced by a ResNet blocks of 3 convolutional layers and a batch normalization layer each, as shown in the right panel of Figure 3. For one model, we add a dropout layer with a rate of 0.5 before the dense layers. As an example, the schematic diagram of the architecture of the two-detector time-series ResNet model is shown in Figure 3. It has 3,862,756 parameters in total. Finally, we point out that the exact hyperparameters we use for our models are unlikely to change the results significantly. We have chosen the specific hyperparameters to give good performance, but a thorough hyperparameter optimization would only have a very small effect on the pipeline’s results.
We build and train our neural network models using TensorFlow (Abadi and others, 2015). We use the Adam optimizer (Kingma and Ba, 2015) with AMSgrad (Reddi et al., 2019). We use $\alpha=0.001$ as our learning rate, $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, and a batch size of 216. We use sparse categorical cross-entropy loss and train for 200 epochs. Our final models are taken from the epoch with the lowest validation loss. We train our models using 4 NVidia A100 GPUs with a data-parallel strategy.
@@ IV.4 Comparison to Other Searches
Now, we will perform a brief comparison of the results of our search to the two most comparable ML-based searches that have been done, with the algorithms AresGW (Koloniari et al., 2025), and AFrame (Marx et al., 2025). Both AresGW and AFrame only searched for BBH events in O3, while we included O1 and O2. With $p_{\textrm{astro}}>0.5$, AresGW detects 40 of the 65 total BBH events in GWTC-3 with two detectors of data and no data quality issues, while AFrame detects 38 of them. Our search detects 49 of the events, although some of those are detected with fairly high False Alarm Rates.
@@ Acknowledgments
This work is supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/).